OpenStudy (anonymous):

solving linear systems of equations; addition -6x=48-6y -4x=50-6y

OpenStudy (bloomlocke367):

hey, I can help :)

OpenStudy (bloomlocke367):

does it matter if you use substitution or elimination to solve?

OpenStudy (anonymous):

it doesnt say. it just wants to solve for x and y .

OpenStudy (bloomlocke367):

okay... well elimination would be easier..

OpenStudy (bloomlocke367):

do you know how to do that?

OpenStudy (anonymous):

no

OpenStudy (bloomlocke367):

actually.. nevermind XD sorry, we'll use substitution..

OpenStudy (bloomlocke367):

pick one equation and decide to solve for either x or y

OpenStudy (anonymous):

ugh i guess the first one x

OpenStudy (bloomlocke367):

okay.. so divide both sides by -6

OpenStudy (bloomlocke367):

what do you get?

OpenStudy (bloomlocke367):

Sorry, I have to go to bed, but @TheSmartOne should be able to help you :)

OpenStudy (bloomlocke367):

where'd they go? sorry @TheSmartOne

TheSmartOne (thesmartone):

It's okay. I'll put out the steps for them for when they return :)

OpenStudy (bloomlocke367):

okay

TheSmartOne (thesmartone):

First of all, Welcome to OpenStudy!! @rhondahyde16 To solve your question, it would be easy to solve by elimination. So if we subtract your two equations as it is, it will look like this: \(\sf ~~~-6x=48-6y\\-\\ ~~~-4x=50-6y\\---------\) If we subtract that, we should be able to eliminate 6y. And to complete that what is \(\sf -6x-(-4x)=?\) And what is \(\sf 48-50=?\) And set them equal to each other like \(\sf -6x-(-4x)=48-50\) And solve for x. Once you get x, plug in the value of x into any equation and solve for y. And you answer will be (x,y)